# A note on the cohomology of moduli spaces of local shtukas

### v1

@@ 3. Cohomology of moduli spaces of local shtukas
on $\mathcal{D}(\mathcal{R}(\mathcal{C}))$, where $f:\mathcal{O}(C)\to\mathcal{Z}(\mathcal{C})$ is the induced homomorphism on centres. Both $\mathcal{O}(\mathcal{C})$ and $\mathcal{Z}(\mathcal{C})$ carry twisting actions of $Y_{ur}$; indeed by choosing a base point they are both isomorphic to quotients of $Y_{ur}$ by finite groups. Since since the Fargues–Scholze construction is compatible with twisting [FS21, Thm. IX.0.5(ii)], $f$ is equivariant for the actions of $Y_{ur}$, and hence finite Galois.
Armed with this, we now consider $\mathcal{D}^{\varphi}(G_{b}(E))$, which is the product of the categories $\mathcal{C}^{\varphi}:=\operatorname{Mod}_{\mathcal{O}(C)/\mathfrak{m}_{\varphi}}\mathcal{C}$, where $\mathcal{C}$ ranges over the Bernstein components of $\mathcal{D}^{C}(G_{b}(E))$ and $\mathfrak{m}_{\varphi}\subseteq\mathcal{O}(C)$ is the maximal ideal corresponding to $\varphi$. By the description of the action of $\operatorname{QCoh}(C)$ on $\mathcal{C}$, $\mathcal{C}^{\varphi}$ is equivalent to $\mathcal{D}(\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C}))$, so we need to show that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is a (classical) semisimple Artinian ring. By [BH03, 8.1, Prop.], $\mathcal{R}(\mathcal{C})$ is an Azumaya algebra over $\mathcal{Z}(\mathcal{C})$. Since $f$ is finite Galois, it follows that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is concentrated in degree 00, where it is a finite product of matrix algebras over $\overline{\mathbb{Q}}_{\ell}$, and in particular semisimple and Artinian, as desired. The rest of the lemma then follows immediately. ∎
@@
In the situation of Theorem 3.5, assume that $\varphi$ is elliptic. Then each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a non-zero finite direct sum of shifted supercuspidal representations. If $\delta$ is one-dimensional, then $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a single supercuspidal representation, up to shift.
@@
We note that Corollary 3.8 has previously appeared in the literature; see [FS21, §X.2] for the case when $S_{\varphi}$ is finite and [Ham22, Cor. 3.11], [BMHN22, Thm. 2.27] for the general case, where it plays a key role in comparing the Fargues–Scholze construction with the local Langlands correspondence. Note that, in light of Corollary 3.8, the Kottwitz conjecture (with respect to the Fargues–Scholze construction) amounts to the image of $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ in the Grothendieck group of $G(E)$ being the supercuspidal representation in the $L$-packet of $\varphi$ corresponding to $\delta$, in the parametrization that uses $\rho$ as a base point. The vanishing conjecture, on the other hand, says that $i_{1}^{\ast}\operatorname{Act}_{\delta}(i_{b\ast}\rho)$ is concentrated in degree 00. In particular, Corollary 3.8 gives a complete understanding of the action of the Weil group. This is in contrast to local approaches to the cohomology of local Shimura varieties predating [FS21], which could say very little about the Weil group action.
@@ 4. Some consequences
In this section we indicate some results towards Conjecture 1.1 that can be obtained from Theorem 3.5 using simple tricks or observations, or recent works on Kottwitz conjecture. We keep the notation and assumptions from §3, and assume additionally that $\varphi$ is elliptic. In this case, the set $\operatorname{Irr}(S_{\varphi},\chi_{b})=\{\delta_{1},\dots,\delta_{r}\}$ is finite, and we write $V_{\varphi,i}$ for $V_{\varphi,\delta_{i}}$. We also assume that $E$ has characteristic 00 throughout this section.
@@ 4.1. The Kottwitz conjecture
For the rest of this section we proceed with a weaker assumption on $\varphi$: We assume that all the $\delta_{i}$ are one-dimensional. This is satisfied in many examples. As noted in the introduction, the local Langlands conjecture predicts the existence of an $L$-parameter $\phi$ attached to $\rho$, and an $L$-packet $\Pi_{\phi}(G)$ of $G(E)$-representations attached to $\phi$. Note that $\Pi_{\phi}(G)$ is a multiset. Write
$i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}^{\prime}[n_{i}]$
where $\pi_{i}^{\prime}$ is supercuspidal and $-[n_{i}]$ denotes a shift by an integer $n_{i}$. In [HKW22], the Kottwitz conjecture is proven modulo the action of $W_{E_{\bullet}}$, under further assumptions on $L$-packets. Combining this with Theorem 3.5, we get Corollary 1.3 from the introduction.
@@
Assume the refined Local Langlands conjecture [Kal16, Conjecture G] and let $\phi$ be the $L$-parameter attached to $\rho$. Assume further that the $L$-packet $\Pi_{\phi}(G)$ only consists of supercuspidal representations, and that all $\delta_{i}$ are one-dimensional. Then
@@
By Theorem 3.5, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{i=1}^{r}(-1)^{n_{i}}\pi_{i}^{\prime}\boxtimes V_{\varphi,i}$ in $\mathrm{Groth}(G(E)\times W_{E_{\bullet}})$. On the other hand, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{j=1}^{s}\pi_{j}\cdot\dim V_{\phi,j}$ in $\mathrm{Groth}(G(E))$ by [HKW22, Thm. 1.0.2], where $\{\pi_{1},\dots,\pi_{s}\}=\Pi_{\phi}(G)$ as multisets. By comparing these expressions we see that the $n_{i}$ are even and that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\varphi}(G)$ as sets. In this comparison, we have used that $\sum_{j=1}^{s}\dim V_{\phi,i}=\dim V=\sum_{i=1}^{r}\dim V_{\varphi,i}$ and that, for any $j$, there exists a $\mu_{\bullet}$ such that $V_{\phi,j}\neq 0$, and similarly for $V_{\varphi,i}$. This proves the theorem. ∎
@@
    While [HKW22, Thm. 1.0.2] is only stated for $\mu_{\bullet}=\mu$ a single cocharacter, its proof goes through for general $\mu_{\bullet}$. We note, however, that Theorem 4.2 for general $\mu_{\bullet}$ follows from the case $\mu_{\bullet}=\mu$, as this case suffices to show that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\phi}(G)$, which is the content of Theorem 4.2 in view of Theorem 3.8.
@@
    As noted in the introduction, this is rather close to the Kottwitz conjecture, but falls short in two aspects. The first issue is that we do not know that $\phi=\varphi$ (indeed, it is currently not known that $\varphi$ is elliptic in most cases when we expect it to be). Lacking this our result looks rather amusing: The $G(E)$-representations in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ arise from $\phi$, whereas the $W_{E_{\bullet}}$-action is given in terms of $\varphi$. The second issue is that, even if $\phi=\varphi$, it is not clear to us that the natural parametrizations of $\Pi_{\phi}(G)$ are the same. At present, the only approach available to these questions is through the cohomology of (global) Shimura varieties. It is known that $\phi=\varphi$ for inner forms of $\operatorname{GL}_{n}$ and $\operatorname{SL}_{n}$ for general $E$ [FS21, HKW22], for inner forms of $\operatorname{GSp}_{4}$ and $\operatorname{Sp}_{4}$ when $E$ is unramified over $\mathbb{Q}_{p}$ [Ham22], for some (similitude) unitary groups in an odd number of variables [BMHN22] when $E=\mathbb{Q}_{p}$, and for $\operatorname{SO}_{2n+1}$ and its unique inner form when $E$ is unramified over $\mathbb{Q}_{p}$ (D.H., unpublished). Comparing the parametrizations of the $L$-packet seems to be more subtle, however.
@@
    We also remark that the extra assumptions of Theorem 4.2 (excluding that $\varphi$ is elliptic) are known to hold in many cases. Examples include inner forms of $\operatorname{SL}_{n}$, $\operatorname{Sp}_{2n}$, $\operatorname{SO}_{2n+1}$ and unitary groups. Moreover, [Kal16, Conjecture G] is known for the regular supercuspidal $L$-packets constructed by Kaletha for all $G$ split over a tame extension of $E$ and for $p$ sufficiently large, in [Kal19]; see [Kal19, FKS23]. We refer to the introduction of [HKW22] for more details.

### v2

@@ 3. Cohomology of moduli spaces of local shtukas
on $\mathcal{D}(\mathcal{R}(\mathcal{C}))$, where $f:\mathcal{O}(C)\to\mathcal{Z}(\mathcal{C})$ is the induced homomorphism on centres. Both $\mathcal{O}(C)$ and $\mathcal{Z}(\mathcal{C})$ carry twisting actions of $Y_{ur}$; indeed by choosing a base point they are both isomorphic to quotients of $Y_{ur}$ by finite groups. Since the Fargues–Scholze construction is compatible with twisting [FS21, Thm. IX.0.5(ii)], $f$ is equivariant for the actions of $Y_{ur}$, and hence finite Galois.
Armed with this, we now consider $\mathcal{D}^{\varphi}(G_{b}(E))$, which is the product of the categories $\mathcal{C}^{\varphi}:=\operatorname{Mod}_{\mathcal{O}(C)/\mathfrak{m}_{\varphi}}\mathcal{C}$, where $\mathcal{C}$ ranges over the Bernstein components of $\mathcal{D}^{C}(G_{b}(E))$ and $\mathfrak{m}_{\varphi}\subseteq\mathcal{O}(C)$ is the maximal ideal corresponding to $\varphi$. By the description of the action of $\operatorname{QCoh}(C)$ on $\mathcal{C}$, $\mathcal{C}^{\varphi}$ is equivalent to $\mathcal{D}(\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C}))$, so we need to show that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is a (classical) semisimple Artinian ring. By [BH03, 8.1, Prop.], $\mathcal{R}(\mathcal{C})$ is an Azumaya algebra over $\mathcal{Z}(\mathcal{C})$. Since $f$ is finite Galois, it follows that $\mathcal{O}(C)/\mathfrak{m}_{\varphi}\otimes^{L}_{\mathcal{O}(C),f}\mathcal{R}(\mathcal{C})$ is concentrated in degree 0, where it is a finite product of matrix algebras over $\overline{\mathbb{Q}}_{\ell}$, and in particular semisimple and Artinian, as desired. The rest of the lemma then follows immediately. ∎
@@
In the situation of Theorem 3.8, assume additionally that $\varphi$ is elliptic. Then each $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a non-zero finite direct sum of shifted supercuspidal representations. If $\delta$ is one-dimensional, then $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ is a single supercuspidal representation, up to shift.
@@
We note that Corollary 3.11 has previously appeared in the literature; see [FS21, §X.2] for the case when $S_{\varphi}$ is finite and [Ham22, Cor. 3.11], [MHN24, Thm. 2.27] for the general case, where it plays a key role in comparing the Fargues–Scholze construction with the local Langlands correspondence. Note that, in light of Corollary 3.11, the Kottwitz conjecture (with respect to the Fargues–Scholze construction) amounts to computing the image of $i^{1\ast}\operatorname{Act}_{\delta}(i^{b}_{\ast}\rho)$ in the Grothendieck group of $G(E)$. The vanishing conjecture, on the other hand, says that $i_{1}^{\ast}\operatorname{Act}_{\delta}(i_{b\ast}\rho)$ is concentrated in degree 0. In particular, Corollary 3.11 gives a complete understanding of the action of the Weil group. This is in contrast to local approaches to the cohomology of local Shimura varieties predating [FS21], which could say very little about the Weil group action.
@@ 4. Some consequences
In this section we indicate some results towards Conjecture 1.1 that can be obtained from Theorem 3.8 using simple tricks or observations, or recent works on Kottwitz conjecture. We will use the same notation as in §3, with a few extra additions. For clarity, we gather this here. These assumptions will be in place throughout the whole of §4.
Setup 4.1.
We let $\rho$ be an irreducible admissible representation of $G_{b}(E)$. We will make the following assumptions, and use the following notation:
  • The Fargues–Scholze parameter $\varphi$ of $\rho$ is elliptic (this forces $\rho$ to be supercuspidal, as noted above).
  • Since $\varphi$ is elliptic, the set $\operatorname{Irr}(S_{\varphi},\chi_{b})$ is finite. We let $\delta_{1},\dots,\delta_{r}$ denote its members, where $r=\#\operatorname{Irr}(S_{\varphi},\chi_{b})$.
  • With $V_{\varphi}$ as in Theorem 3.8, write $V_{\varphi,i}=\operatorname{Hom}_{S_{\varphi}}(\delta_{i},V_{\varphi})$, viewed as a $W_{E_{\bullet}}$-representation.
  • Finally, assume that $E$ has characteristic 0.
@@ 4.1. The Kottwitz conjecture
We now return to the general case. In [HKW22, Thm 1.0.2], the Kottwitz conjecture is proven after disregarding the action of $W_{E_{\bullet}}$, under further assumptions on $L$-packets. We wish to combine this with Theorem 3.8 in order to deduce Corollary 1.3 from the introduction. This will use the following assumptions and notation:
Setup 4.3.
Assume the refined local Langlands conjecture in the form of [Kal16, Conjecture G]. In particular, we have an $L$-parameter $\phi$ attached to $\rho$, and an $L$-packet $\Pi_{\phi}(G)$ (a set) of $G(E)$-representations. We assume that $\phi$ is discrete (i.e. that $S_{\phi}/Z(\widehat{G})^{Q}$ is finite), with centralizer $S_{\phi}$, and that all members of $\Pi_{\phi}(G)$ are supercuspidal.
Under these assumptions, [HKW22, Thm. 1.0.2] asserts that
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\pi\in\Pi_{\phi}(G)}\dim\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},V_{\phi})\cdot\pi$
in the Grothendieck group of $G(E)$-representations (recall from the introduction that $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho):=\sum_{n}(-1)^{n}H^{n}(R\Gamma(G,b,\mu_{\bullet})[\rho]$). Here $\delta_{\pi,\rho}$ is a certain (not necessarily irreducible) representation of $S_{\phi}$ attached to the pair $(\pi,\rho)$ with $Z(\widehat{G})^{Q}$ acting by $\chi_{b}$, described in [HKW22, §2.3]. We will need the following lemma concerning the representations $\delta_{\pi,\rho}$.
Lemma 4.4.
Any $\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})$ occurs as a subrepresentation of $\bigoplus_{\pi\in\Pi_{\phi}(G)}\delta_{\pi,\rho}$. If all $\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})$ are one-dimensional, then we have
$\sum_{\pi\in\Pi_{\phi}(G)}\dim\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},W)=\dim\operatorname{Hom}_{S_{\phi}}(\oplus_{\pi\in\Pi_{\phi}(G)}\delta_{\pi,\rho},W)\geq\dim W$
for any $W\in\operatorname{Rep}(S_{\phi})$ with $Z(\widehat{G})^{Q}$ acting as $\chi_{b}$.
Proof.
For this lemma, we will refer to [HKW22, §2.3] and the references within for unexplained terminology, notation, and facts (see also [Kal16, §4.5]). There is a group $S_{\phi}^{+}$ arising as a covering group $f:S_{\phi}^{+}\to S_{\phi}$. Set $Z(\widehat{\bar{G}})^{+}:=f^{-1}(Z(\widehat{G})^{Q})$; this group is central in $S_{\phi}^{+}$. It induces a map $\pi_{0}(Z(\widehat{\bar{G}}))\to\pi_{0}(S_{\phi}^{+})$. If $\lambda$ is a character of $)\pi_{0}(Z(\widehat{\bar{G}}$)), we will say that a representation of $\pi_{0}(S_{\phi}^{+})$ has character $\lambda$ if $)\pi_{0}(Z(\widehat{\bar{G}}$)) acts as $\lambda$ on it. According to [Kal16, Conjecture G], there is a bijection $\pi\to\tau_{z,\mathfrak{m},\pi}$ between $\Pi_{\phi}(G)$ and irreducible representations of $\pi_{0}(S_{\phi}^{+})$ with a certain character $\lambda_{z}$, and there is an irreducible representation $\tau_{z,\mathfrak{m},\rho}$ of $\pi_{0}(S_{\varphi}^{+})$ associated with $\rho$, which has a certain character $\lambda_{z}+\lambda_{z_{b}}$. By definition, $\delta_{\pi,\rho}$ is the descent of $\tau_{z,\mathfrak{m},\pi}^{\ast}\otimes\tau_{z,\mathfrak{m},\rho}$ to $S_{\phi}$ along $f$ (and hence $\lambda_{z_{b}}$ corresponds to $\chi_{b}$). With this in mind, let $\delta$ be an irreducible representation of $S_{\phi}$ with character $\chi_{b}$. Pullback along $f$ gives an equivalence of representations of $S_{\phi}$ with character $\chi_{b}$ and representations of $\pi_{=}(S_{\phi}^{+})$ with character $\lambda_{z_{b}}$, so we also let $\delta$ denote the corresponding representation of $\pi_{0}(S_{\phi}^{+})$. Then $\delta^{\ast}\otimes\tau_{z,\mathfrak{m},\rho}$ has character $\lambda_{z}$, and can therefore be written as a sum of the $\tau_{z,\mathfrak{m},\pi}$ (possibly with multiplicities). Moreover, $\delta$ is a summand of $(\delta^{\ast}\otimes\tau_{z,\mathfrak{m},\rho})^{\ast}\otimes\tau_{z,\mathfrak{m},\rho}$. This proves the first part of the lemma. To deduce the second part, we now have $\oplus_{\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})}\delta\subseteq\oplus_{\pi\in\Pi_{\phi}(G)}\delta_{\pi,\rho}$ (non-canonically), so
$\dim\operatorname{Hom}_{S_{\phi}}(\oplus_{\pi\in\Pi_{\phi}(G)}\delta_{\pi,\rho},W)\geq\dim\operatorname{Hom}_{S_{\phi}}(\oplus_{\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})}\delta,W)\geq\dim W,$
where the second inequality uses that the $\delta$ are one-dimensional. ∎
With these preparations, we can deduce Corollary 1.3.
@@
Assume Setups 4.1 and 4.3, and that all $\delta_{i}$ and all $\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})$ are one-dimensional. Write $i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}^{\prime}[n_{i}]$, where $\pi_{i}^{\prime}$ is supercuspidal and $n_{i}]$ denotes a shift by an integer $n_{i}$ (we can do this by Corollary 3.11). Then
@@
By Theorem 3.8, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{i=1}^{r}(-1)^{n_{i}}\pi_{i}^{\prime}\boxtimes V_{\varphi,i}$ in $\mathrm{Groth}(G(E)\times W_{E_{\bullet}})$. In particular, $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{i=1}^{r}(-1)^{n_{i}}\dim V_{\varphi,i}\cdot\pi_{i}^{\prime}$ in $\mathrm{Groth}(G(E))$, where we have that
$\sum_{i=1}^{r}(-1)^{n_{i}}\dim V_{\varphi,i}\leq\sum_{i=1}^{r}\dim V_{\varphi,i}=\dim V$
since the $\delta_{i}$ run through $\operatorname{Irr}(S_{\varphi},\chi_{b})$. On the other hand,
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\pi\in\Pi_{\phi}(G)}\dim\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},V_{\phi})\cdot\pi$
in $\mathrm{Groth}(G(E))$ by [HKW22, Thm. 1.0.2]. By Lemma 4.4, we have $\sum_{\pi\in\Pi_{\phi}(G)}\dim\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},V_{\phi})\geq\dim V$. Note that, for a fixed $i$ (or $\pi$) and varying $\mu_{\bullet}$, some $V_{\varphi,i}$ (or $\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},V_{\phi})$) will be non-zero. Using this fact and comparing the two inequalities we have derived so far, we see that the $n_{i}$ are even and that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\varphi}(G)$, as desired. ∎
@@
    While [HKW22, Thm. 1.0.2] is only stated for $\mu_{\bullet}=\mu$ a single cocharacter, its proof goes through for general $\mu_{\bullet}$. We note, however, that Theorem 4.5 for general $\mu_{\bullet}$ follows from the case $\mu_{\bullet}=\mu$ (in light of Corollary 3.11), as this case suffices to show that $\{\pi_{1}^{\prime},\dots,\pi_{r}^{\prime}\}=\Pi_{\phi}(G)$ and that the $n_{i}$ are even.
@@
    As noted in the introduction, this is rather close to the Kottwitz conjecture, but falls short in two aspects. Firstly, we do not know that $\phi=\varphi$ (indeed, we do not known that $\varphi$ is elliptic in most cases it is expected to be). Lacking this, our result looks rather amusing: The $G(E)$-representations in $R\Gamma(G,b,\mu_{\bullet})[\rho]$ arise from $\phi$, whereas the $W_{E_{\bullet}}$-action is given in terms of $\varphi$. Moreover, even if $\phi=\varphi$, it is not clear to us how the resulting parametrizations of $\Pi_{\phi}(G)$ compare. At present, the only approach available to these questions is through the cohomology of (global) Shimura varieties. It is known that $\phi=\varphi$ for inner forms of $\operatorname{GL}_{n}$ and $\operatorname{SL}_{n}$ for general $E$ [FS21, HKW22], for inner forms of $\operatorname{GSp}_{4}$ and $\operatorname{Sp}_{4}$ when $E$ is unramified over $\mathbb{Q}_{p}$ [Ham22], for some (similitude) unitary groups in an odd number of variables [MHN24] when $E=\mathbb{Q}_{p}$, and for special orthogonal and unramified unitary groups when $E$ is unramified [Pen25]. Comparing the parametrizations of the $L$-packet seems to be more subtle, however.
@@
    One-dimensionality is known to hold in many cases. Examples include inner forms of $\operatorname{SL}_{n}$, $\operatorname{Sp}_{2n}$, $\operatorname{SO}_{2n+1}$ and unitary groups. Moreover, [Kal16, Conjecture G] is known for the regular supercuspidal $L$-packets constructed by Kaletha for all $G$ split over a tame extension of $E$ and for $p$ sufficiently large, in [Kal19]; see [Kal19, FKS23]. We refer to the introduction of [HKW22] for more details.